the length of the second side of a trianglei s 2 inches less than the lenght of the first side. the length of the third side is 12 inches more than the length of the first side. the perimeter is 73 inches. find the lenth of each side of the triangle.
Answer by checkley71(8403)

PERIMETER OF A TRIANGLE=SIDE1+SIDE2+SIDE3
LET X=THE FIRST SIDE.
THE SECOND SIDE=X-2
THE THIRD SIDE=X+12
73=X+X-2+X+12
73=3X+10
3X=73-10
3X=63
X=63/3
X=21 FOR THE FIRST SIDE.
21-2=19 FOR THE SECOND.
21+12=33 FOR THE THIRD.
PROOF:
21+19+33=73
73=73